OpenStudy (anonymous):

Interest Problem: Julian deposits money into two different savings accounts. He deposits $600 more into Account B than Account A. After a certain amount of time, Account A has earned $180 at a rate of 4%, and Account B has earned $375 at a rate of 5%. How much did Julian initially deposit into each account? a. Use I=prt where I is the earned interest, p is the principal, r is the interest rate and t is the time. b. Identify the given information. What do you know about P? I c. Enter the given information into the equation. d. Solve the equation. e. State the solution.

OpenStudy (anonymous):

ok i will try is it multiple choice?

OpenStudy (anonymous):

nope. finally someone at least looks at it

OpenStudy (anonymous):

abcd and all that are the steps

OpenStudy (anonymous):

ok i got this just hang on

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

what is the name of your math book?

OpenStudy (anonymous):

i don't have one its florida virtual school

OpenStudy (anonymous):

ahh ok i am homeschooled as well

OpenStudy (anonymous):

no I'm not home schooled its just an extra class cuz i like to get ahead of my grade

OpenStudy (anonymous):

Account B = Account A +600, this information should help you set up (a)

OpenStudy (anonymous):

i guess i will let this guy handle things

OpenStudy (anonymous):

i saw something like that on yahoo answers but i need to use I= prt

OpenStudy (anonymous):

idk if he's right

OpenStudy (anonymous):

ok hang on

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

new viewer in the house woot woot

OpenStudy (anonymous):

im sorry cant figure it out

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

come on vadatajs you can do it!

OpenStudy (anonymous):

I=prt for A you have $180=(x)(4)(t), for B you have $375=(x+600)(5)(t)

OpenStudy (anonymous):

im so sorry im just not that smart

OpenStudy (anonymous):

where x is the principal, and t is the time, since neither are stated, you must use variables

OpenStudy (anonymous):

this is algebra 1

OpenStudy (anonymous):

i know i just started it

OpenStudy (anonymous):

@Vadatajs can u solve it please

OpenStudy (anonymous):

oh yeah this is at the end. good luck. i hate taking online classes -_-

OpenStudy (deoxna):

I=pr^t For account A you have $180=(x)(1.04)^(t), and for B you have $375=(x+600)(1.05)^(t). Am I right? Im pretty sure the t should be exponential.

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

To solve a problem involving simple interest and two accounts, use the equation I/pr= I/pr where I represents the amount of interest earned, P represents the initial amount deposited, and r represents the interest rate. Make sure the left side of the equation contains the information for one account, and the right side contains the information for the other account. Enter the given information into the equation and solve for the unknown variable.

OpenStudy (deoxna):

Ok its not exponential ( http://www.teacherschoice.com.au/maths_library/money/simple_interest.htm), that makes it easier... So the new equations are Account A is $180=(x)(.04)(t) Account B is $375=(x+600)(1.05)(t) 4500/t=x 375=4725t/t + 630t 4325=630t t=6.904 x=651.72

OpenStudy (anonymous):

are you positive thats right?

OpenStudy (anonymous):

@ZhangYan is it right?

OpenStudy (anonymous):

let A = amount deposited in A B = A+600 A*0.04 =180; A= 180/0.04 = 4,500 B*0.05=375; B = 375/0.05 = 7,500 principal*interest rate = interest earns principal = A or B interest rate is presumed to be EAR and not APR, which would introduce timing problems and would have explained the discrepancies above.

OpenStudy (anonymous):

hope this helps :)

OpenStudy (anonymous):

so can you work it out for me please i have no idea

OpenStudy (deoxna):

Sorry about that, the answer is t=5, x=900

OpenStudy (anonymous):

i already did. I'll show u the example Let the deposit in Account A be x 4% of x = 180 Assuming the deposit is for one year, x = 4,500 Deposit in Account B = x + 600 = 5100 Interest at 5% for one year = 255 Interest earned as per the problem is 375 = 255 x 1.47 The deposit in Account B is $7,500 The deposit in Account A is $6,615

OpenStudy (anonymous):

is the work correct though @DeoxNA

OpenStudy (deoxna):

Account A is $180=(x)(.04)(t) Account B is $375=(x+600)(1.05)(t) 4500/t=x 375=0.05xt+30t 375=225t/t + 30t t=5 x=900

OpenStudy (anonymous):

and thanks @ZhangYan but i needed it in the right equation of i = prt

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

wait but @DeoxNA u didn't use principal

OpenStudy (anonymous):

im confused af

OpenStudy (deoxna):

The principal is x

OpenStudy (anonymous):

oh okay and whats r the numbers?

OpenStudy (anonymous):

I'm pretty sure its wrong cuz theres only supposed to be one solution dn u gave two...

OpenStudy (anonymous):

yay viewers!

OpenStudy (anonymous):

are u guys just gonna stare at the question or answer it

OpenStudy (anonymous):

A = P(1 + i)^n A = Amount P = Principal invested i = interest rate n = number of years Let x = amount deposited in account A Amount deposited in account B = x + 600 1) (x + 180) = x(1.04)^n 2) (x + 600 + 375) = (x + 600)(1.05)^n >>> ( x+975) = (x + 600)*1.05^n

OpenStudy (anonymous):

does this help this time

OpenStudy (anonymous):

makes since to me thanks

OpenStudy (anonymous):

@nicole1217 are you there?

OpenStudy (anonymous):

ok then leave the rest to me Bye

OpenStudy (anonymous):

I'm sorry i don't think thats right

OpenStudy (anonymous):

yes it is correct...

OpenStudy (anonymous):

are you sure I'm just confused about the format

OpenStudy (anonymous):

why it doesn't make any sense to you well @countonme123 understood....

OpenStudy (deoxna):

@nicole1217 I'm pretty sure the answer I gave you right, the question was asking for the initla money deposited, which is the variable x in my equation. The problem requieres two variables to be solved, but the answer to the question is x=900.

OpenStudy (anonymous):

everyone doesnt learn the same